Light of wavelength 650 nm falls on a slit that is 3.90×10−3 mm wide.

How far the first bright diffraction fringe is from the strong central maximum if the screen is 13.0 m away.

This is my last attempt; I've put 3.25 m, 2.20 m, 2.17 m
Answer #1

Wavelength of light = 1 = 650 nm Inm=10 9 m » = 650 X 10 m = 6-58107m width of slit = Q = 3.90 X10-3 mm : I mm = 10-3 m ) a =- y, = tan Q = y = (13.0m) tan (14.47750) =) Jy = (13.0 m) 10.2582) => Y, = 3.3566 m Ž 3.36 m So, first bright diffraction fr